OpenStudy (anonymous):

Express the concentration of a 0.0100 M aqueous solution of fluoride, F–, in mass percentage and in parts per million. Assume the density of the solution is 1.00 g/mL.

OpenStudy (aaronq):

Find the mass of the solution and of the solute. Useful formulas: \(Molarity=\dfrac{n_{solute}}{L_{solution}}\) \(moles=\dfrac{mass}{Molar~mass}\) \(density=\dfrac{mass}{Volume}\) for the second part: 1 mg/mL = 1 ppm

OpenStudy (wolfe8):

And welcome to OpenStudy!